# Handling Reward Misspecification in the Presence of Expectation Mismatch

### v1

@@ 4 Identifying Expectation-Aligned Policies
This follows directly from the definitions of the expectation set and the planning function.
$\begin{split}\max_{x}\sum_{s,a}x(s,a)\times r(s)\\[-8.0pt] \text{s.t.}\quad\forall s\in S,\sum_{a}x(s,a)=(\delta(s,s_{0})+\\[-5.0pt] ~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\gamma\times\sum_{s^{\prime},a^{\prime}}x(s^{\prime},a^{\prime})\times T^{H}(s,a,s^{\prime}))\\[-8.0pt] \sum_{a}x(s_{0},a)=V^{*}_{s_{0}}\\[-8.0pt] \sum_{a}x(s_{i},a)=0\end{split}$ (3)
This brings us to the proposition
Proposition 4.
Now with both sets $\widehat{\mathcal{S}}^{\mathcal{F}}$ and $\widehat{\mathcal{S}}^{\mathcal{G}}$ calculated, the objective of the robot is to generate a policy that covers all of the states in $\widehat{\mathcal{S}}^{\mathcal{G}}$, while avoiding the states in $\widehat{\mathcal{S}}^{\mathcal{F}}$. This may not always be possible, so we will need the system to query the user about whether a state can be visited or not. Based on their response, we can update the sets and try to see if we can generate a policy that satisfies the requirement. Now the basic formulation takes the form of the form of an LP:
$\begin{split}\max_{x,\mathbb{D}_{F},\mathbb{D}_{G}}-1\times\sum_{d\in\mathbb{D}_{G}\cup\mathbb{D}_{F}}d\\ \text{s.t.}\quad\forall s\in S,\sum_{a}x(s,a)=\delta(s,s_{0})+\\[-8.0pt] \gamma\times\sum_{s^{\prime},a^{\prime}}x(s^{\prime},a^{\prime})\times T^{R}(s,a,s^{\prime})\\[-8.0pt] \forall d\in\mathbb{D}_{G}\cup\mathbb{D}_{F},d\geq 0\\[-3.0pt] \forall s_{i}\in\widehat{\mathcal{S}}^{\mathcal{F}}\sum_{a}x(s_{i},a)-d_{i}=0\\[-8.0pt] \forall s_{j}\in\widehat{\mathcal{S}}^{\mathcal{G}}\sum_{a}x(s_{j},a)+d_{j}\geq 0\end{split}$ (4)
Theorem 2.
Proof Sketch.
When such solutions are not found, the $d$ variables that take positive values tell us potential states we can query. After the query if the user says a state shouldn’t be visited or should be visited the corresponding states are removed from $\widehat{\mathcal{S}}^{\mathcal{F}}$ or $\widehat{\mathcal{S}}^{\mathcal{G}}$ and move them to the set $\mathcal{S}^{\mathcal{F}}_{*}$ and $\mathcal{S}^{\mathcal{G}}_{*}$. Once these sets become non-empty, we use an updated form of Equation 4, which now has hard constraints of the form.
$\begin{split}\forall s_{i}\in\mathcal{S}^{\mathcal{F}}_{*},\sum_{a}x(s_{i},a)=0\\[-8.0pt] \forall s_{i}\in\mathcal{S}^{\mathcal{G}}_{*},\sum_{a}x(s_{i},a)\geq 0\end{split}$ (5)
Theorem 2 also holds for this new LP. This process is repeated until a policy is found or if there are no more elements in $\widehat{\mathcal{S}}^{\mathcal{F}}\cup\widehat{\mathcal{S}}^{\mathcal{G}}$. We provide a pseudo-code for the entire process in Algorithm 1. The algorithm is guaranteed to terminate for any finite state MDP.
@@
     $X,\mathbb{D}_{G},\mathbb{D}_{F}\leftarrow Solve(LP)$
@@
     $\mathcal{S}^{\mathcal{G}}_{*}\leftarrow\mathcal{S}^{\mathcal{G}}_{*}\cup QueryGoal(\mathbb{S}^{\prime}_{\mathcal{G}})$
     $\mathcal{S}^{\mathcal{F}}_{*}\leftarrow\mathcal{S}^{\mathcal{F}}_{*}\cup QueryForbidden(\mathbb{S}^{\prime}_{\mathcal{F}})$
@@ 5 Related Works
Given the ever-increasing capabilities of AI agents, there has been a lot of work that has looked at how to handle partially specified or misspecified objectives. Most work in this area tends to be motivated by the need to ensure safety. This is fueled by the recognition that people tend to be bad at correctly specifying their objectives Booth et al. (2023); Milli et al. (2017), and optimizing for incorrectly specified objectives could have disastrous results Hadfield-Menell et al. (2016). One could roughly categorize the work done to develop more general methods to address this problem into two broad categories.
In the first, the reward function is mostly taken to be true, but they assume there is a set of hard constraints, usually partially specified, on what state features could be changed Weld and Etzioni (1994), which are usually referred to as side-effects. As such, these works are usually framed in terms of avoiding negative side effects. Many works in this direction focus on the problem of how to identify potential side-effect features (cf. Zhang et al. (2018); Saisubramanian et al. (2022); Mahmud et al. (2023a)), mostly by querying the users. Recent work has also looked at extending these methods to non-markovian side-effects Srivastava et al. (2023) and even to possible epistemic side effects Klassen et al. (2023). In addition to querying, there are also works that look at minimizing potential side effects when coming up with behavior Klassen et al. (2022a). Previous work includes work in the context of factored planning problems (cf. Klassen et al. (2022b); Mechergui and Sreedharan (2023)) and reinforcement learning Vamplew et al. (2021).
The second category treats the reward function provided by the user as merely constituting a partial specification or an observation of the true reward function held by the user. A canonical example is inverse reward design Hadfield-Menell et al. (2017), which uses the specified reward function as a proxy to infer the true objective. For a good balance between conservatism and informativeness, Krasheninnikov et al. Krasheninnikov et al. (2021) proposes an algorithm to infer the reward function from two conflicting reward function sources. Mahmud et al. Mahmud et al. (2023b) use explanations within their framework to verify and improve reward alignment. Pan et al. Pan et al. (2022) proposes anomaly detection to tackle reward hacking due to misspecification. Connected to this stream of work is the CIRL framework Hadfield-Menell et al. (2016), which eschews an agent from having its own reward function, but tries to maximize the human reward function. However, the agent estimate of this reward function may also be incorrect or incomplete.
@@
Another stream of work that is relevant to the current problem is that of imitation learning Torabi et al. (2019) or learning from demonstration Argall et al. (2009). Under this regime, the teacher demonstrates a particular course of action to an agent, and it is expected to either imitate or generate behaviors that match the intent behind the demonstration. Apprenticeship learning Abbeel and Ng (2004), and other methods that use inverse reinforcement learning (cf. Arora and Doshi (2021)) treat the reward function as an unknown entity and use the demonstration and any knowledge about the demonstrator to identify the true reward function. The objective then becomes to identify a policy that maximizes this learned reward function. In some sense, our work inverts this paradigm and starts from a specified reward function and tries to recreate what expectations may have led to this reward function. In addition to all the work on partially specified rewards, our work is also connected to preference/model elicitation (cf. Boutilier (2002); Chen and Pu (2004)). Our proposed method can be thought of as a form of preference elicitation, except we focus on learning some specific kinds of preferences. It is also worth noting that there have been some efforts within the inverse-reinforcement learning community to formalize misspecification (cf. Skalse and Abate (2023)). However, we see such work being complementary to our effort.
@@
Table 1: The table reports the number of queries raised and the time taken by our method. For IRD, it shows the number of expectations violated by the generated policy and the time taken.
6 Evaluation
Our primary goal with the empirical evaluation was to compare our proposed method against two baseline methods selected from the two groups of works described in Section 5. Specifically, we wanted to test,
”How the method described in Section 4 compared with existing methods in terms of (a) computational efficiency (measured in terms of time-taken), (b) overhead placed on the user (number of queries) and (c) ability to satisfy user-expectations.”
In particular, we selected Minimax-Regret Querying (MMRQ-k)Zhang et al. (2018) (with query size $k=2$ and all features set to unknown) and a modified form of the Inverse Reward Design method Hadfield-Menell et al. (2017). For the modified IRD, we avoided the expensive step of calculating posterior distribution over the reward function set and instead directly used $\widehat{S}^{G}$ and $\widehat{S}^{F}$. For the expected reward function, a high positive reward was assigned for $\widehat{S}^{G}$ and a negative one for $\widehat{S}^{F}$. We then try to solve the MDP for this reward function using an LP planner.
Test Domains.
We tested our method and baseline on five domains. Most of these are standard benchmark tasks taken from the SimpleRL libraryAbel (2019). Since all examples were variations of a basic grid-world domain, we considered four different sizes and five random instantiations of each grid size (obtained by randomizing the initial state, goal state, and location of objects). For each of the tasks, the expectation set consists of reaching the goal state and avoiding some random states in the environment. The human models were generated by modifying the original task slightly.
The walkway involves a simple grid world where the robot can use a movable walkway to reach certain states easily, but the human is unaware of it. Obstacles involve the robot navigating to a goal while avoiding obstacles (the human model includes incorrect information about the obstacles). Four rooms Sutton et al. (1999) involves the robot navigating through connected rooms, but in the human model, the use of certain doors may not be allowed. In Puddle, the robot needs to travel to the goal while avoiding certain terrain types, while the human model may be wrong about the location of various terrain elements. Finally, in maze, the robot needs to navigate a maze, while the human model may have incorrect information about what paths are available.
Evaluation.
All the baselines were run with a time-bound of 30 minutes per problem. All experiments were run on AlmaLinux 8.9 machine with 32GB RAM and 16 Intel(R) Xeon(R) 2.60GHz CPUs. We used CPLEX Bliek1ú et al. (2014) as our LP solver. First,we found that the MMRQ-k method could not solve any of the instances. This was because the runtime was dominated by the time needed to calculate the exponential number of dominant policies. This shows how computationally expensive methods from the first group are. On the other hand, our method is much faster, and even in the largest grids, it took less than five minutes (Table 1) and only required very few queries. Note that the maximum number of queries that could be raised in each case corresponds to the total state space. In each case considered here the number of queries raised was substantially smaller than the state space size. Thus showing the effectiveness of our method compared to existing query methods.

### v2

@@ 4 Identifying Expectation-Aligned Policies
This follows directly from the definitions of the expectation set and the planning function (a longer sketch is provided in Appendix 8.1).
@@
     $Solve,X,\mathbb{D}_{G},\mathbb{D}_{F}\leftarrow Solve(LP)$
@@
     $\mathcal{S}^{\mathcal{G}}_{*}\leftarrow\mathcal{S}^{\mathcal{G}}_{*}\cup QueryGoal(\mathbb{S}^{\prime}_{\mathcal{G}})$, $\mathcal{S}^{\mathcal{F}}_{*}\leftarrow\mathcal{S}^{\mathcal{F}}_{*}\cup QueryForbidden(\mathbb{S}^{\prime}_{\mathcal{F}})$
@@
$\begin{split}\max_{x}\sum_{s,a}x(s,a)\times r(s,a)\\[-8.0pt] \text{s.t.}\quad\forall s\in S,\sum_{a}x(s,a)=\delta(s,s_{0})+\gamma\times\sum_{s^{\prime},a^{\prime}}x(s^{\prime},a^{\prime})\times T^{H}(s^{\prime},a^{\prime},s)\\[-8.0pt] \sum_{a}x(s_{0},a)\times r(s,a)=V^{*}_{s_{0}};~{}~{}~{}~{}\sum_{a}x(s_{i},a)=0\end{split}$ (3)
This brings us to the proposition.
Proposition 4.
Now with both sets $\widehat{\mathcal{S}}^{\mathcal{F}}$ and $\widehat{\mathcal{S}}^{\mathcal{G}}$ calculated, the objective of the robot is to generate a policy that covers all of the states in $\widehat{\mathcal{S}}^{\mathcal{G}}$, while avoiding the states in $\widehat{\mathcal{S}}^{\mathcal{F}}$. This may not always be possible, so we will need the system to query the user about whether a state $s\in\widehat{\mathcal{S}}^{\mathcal{F}}$ is actually part of ${\mathcal{S}}^{\mathcal{F}}$ (i.e., “do I need to avoid $s$?"), and if a state $s^{\prime}\in\widehat{\mathcal{S}}^{\mathcal{G}}$ is actually part of ${\mathcal{S}}^{\mathcal{G}}$ (“do I need to visit $s^{\prime}$?"). Based on their response, we can update the sets and try to see if we can now generate a policy that avoids remaining states in the updated set of forbidden states and visits all the states in the updated goal state set. We can identify whether such a policy exists and possible states to query by using the LP:
$\begin{split}\max_{x,\mathbb{D}_{F},\mathbb{D}_{G}}-1\times\sum_{d\in\mathbb{D}_{G}\cup\mathbb{D}_{F}}d\\ \text{s.t.}\quad\forall s\in S,\sum_{a}x(s,a)=\delta(s,s_{0})+\gamma\times\sum_{s^{\prime},a^{\prime}}x(s^{\prime},a^{\prime})\times T^{R}(s^{\prime},a^{\prime},s)\\[-8.0pt] \forall d\in\mathbb{D}_{G}\cup\mathbb{D}_{F},d\geq 0\\ \forall s_{i}\in\widehat{\mathcal{S}}^{\mathcal{F}}\sum_{a}x(s_{i},a)-d_{i}=0;~{}~{}~{}\forall s_{j}\in\widehat{\mathcal{S}}^{\mathcal{G}}\sum_{a}x(s_{j},a)+d_{j}\geq 0\end{split}$ (4)
Theorem 2.
Proof Sketch.
When such solutions are not found, the $d$ variables that take positive values tell us potential states we can query. After the query if the user says a state shouldn’t be visited or should be visited the corresponding states are removed from $\widehat{\mathcal{S}}^{\mathcal{F}}$ or $\widehat{\mathcal{S}}^{\mathcal{G}}$ and moved to the set $\mathcal{S}^{\mathcal{F}}_{*}$ (set of known forbidden states) and $\mathcal{S}^{\mathcal{G}}_{*}$ (set of known goal states). Once these sets become non-empty, we use an updated form of Equation 4, which now has hard constraints of the form.
$\begin{split}\forall s_{i}\in\mathcal{S}^{\mathcal{F}}_{*},\sum_{a}x(s_{i},a)=0;~{}\forall s_{i}\in\mathcal{S}^{\mathcal{G}}_{*},\sum_{a}x(s_{i},a)\geq 0\end{split}$ (5)
Theorem 2 also holds for this new LP. If the user answers that a state is neither a forbidden nor a goal state, then the state is simply removed from the corresponding superset. This process is repeated until a policy is found, if there are no more elements in $\widehat{\mathcal{S}}^{\mathcal{F}}\cup\widehat{\mathcal{S}}^{\mathcal{G}}$, or if the LP is unsolvable.
Algorithm 1 provides the pseudo-code for the query procedure. The procedure $CreateLP$ creates an LP of the form described in Equation 4, for the current estimates for $\widehat{\mathcal{S}}^{\mathcal{F}}$ and $\widehat{\mathcal{S}}^{\mathcal{G}}$ (along with known ones). The estimates are refined through queries with the user (represented by the procedures $QueryGoal$ and $QueryForbidden$). The algorithm stops if a solution is found that doesn’t require any additional queries, if the LP is unsolvable, or if it runs out of states to query.
Proposition 5.
Algorithm 1, is guaranteed to exit in finite steps for all finite state space MDPs.
The above proposition holds since, in the worst case, the algorithm would query about every state in the MDP. Appendix 8.2 provides an extension of the method for a planning function based on a noisy rational model.
@@ 5 Related Works
Given the ever-increasing capabilities of AI agents, there has been a lot of work that has looked at how to handle partially specified or misspecified objectives. Most work in this area tends to be motivated by the need to ensure safety. This is fueled by the recognition that people tend to be bad at correctly specifying their objectives [Booth et al., 2023, Milli et al., 2017], and optimizing for incorrectly specified objectives could have disastrous results [Hadfield-Menell et al., 2016]. One could roughly categorize the work done to develop more general methods to address this problem into two broad categories.
In the first, the reward function is mostly taken to be true, but they assume there is a set of hard constraints, usually partially specified, on what state features could be changed [Weld and Etzioni, 1994], which are usually referred to as side-effects. As such, these works are usually framed in terms of avoiding negative side effects. Many works in this direction focus on the problem of how to identify potential side-effect features (cf. [Zhang et al., 2018, Saisubramanian et al., 2022, Mahmud et al., 2023b]), mostly by querying the users. Recent work has also looked at extending these methods to non-markovian side-effects [Srivastava et al., 2023] and even to possible epistemic side effects [Klassen et al., 2023]. In addition to querying, there are also works that look at minimizing potential side effects when coming up with behavior [Klassen et al., 2022b]. Previous work includes work in the context of factored planning problems (cf. [Klassen et al., 2022a, Mechergui and Sreedharan, 2023]) and reinforcement learning [Vamplew et al., 2021].
The second category treats the reward function provided by the user as merely constituting a partial specification or an observation of the true reward function held by the user. A canonical example is inverse reward design [Hadfield-Menell et al., 2017], which uses the specified reward function as a proxy to infer the true objective. For a good balance between conservatism and informativeness, Krasheninnikov et al. [2021] proposes an algorithm to infer the reward function from two conflicting reward function sources. Mahmud et al. [2023a] use explanations within their framework to verify and improve reward alignment. Pan et al. [2022] proposes anomaly detection to tackle reward hacking due to misspecification. This is also related to the CIRL framework [Hadfield-Menell et al., 2016], which eschews an agent from having its own reward function, but tries to maximize the human reward function. However, the agent estimate of this reward function may also be incorrect or incomplete.
@@
Another stream of work that is relevant to the current problem is that of imitation learning [Torabi et al., 2019] or learning from demonstration [Argall et al., 2009]. Under this regime, the teacher demonstrates a particular course of action to an agent, and it is expected to either imitate or generate behaviors that match the intent behind the demonstration. Apprenticeship learning [Abbeel and Ng, 2004], and other methods that use inverse reinforcement learning (cf. [Arora and Doshi, 2021]) treat the reward function as an unknown entity and use the demonstration and any knowledge about the demonstrator to identify the true reward function. The objective then becomes to identify a policy that maximizes this learned reward function. In some sense, our work inverts this paradigm and starts from a specified reward function and tries to recreate what expectations may have led to this reward function. In addition to all the work on partially specified rewards, our work is also connected to preference/model elicitation (cf. [Boutilier, 2002] and [Chen and Pu, 2004]). Our proposed method can be thought of as a form of preference elicitation, except we focus on learning some specific kinds of preferences. It is also worth noting that there have been some efforts within the inverse-reinforcement learning community to formalize misspecification (cf. [Skalse and Abate, 2023]). However, we see such work being complementary to our effort.
6 Evaluation
Our primary goal with the empirical evaluation was to compare our proposed method against two baseline methods selected from the two groups of works described in Section 5. Specifically, we wanted to test:
"How the method described in Section 4 compared with existing methods in terms of (a) computational efficiency (measured in terms of time-taken), (b) overhead placed on the user (number of queries) and (c) ability to satisfy user-expectations."
In particular, we selected Minimax-Regret Querying (MMRQ-k) [Zhang et al., 2018] (with query size $k=2$ and all features set to unknown) and a modified form of the Inverse Reward Design method [Hadfield-Menell et al., 2017]. To simplify the setting, we considered an MDP formulation where rewards were associated with just states as opposed to states and actions. For the modified IRD, we avoided the expensive step of calculating posterior distribution over the reward function set and instead directly used $\widehat{S}^{G}$ and $\widehat{S}^{F}$. For the expected reward function, a high positive reward was assigned for $\widehat{S}^{G}$ and a negative one for $\widehat{S}^{F}$. We then try to solve the MDP for this reward function using an LP planner. For the query-based methods, we simply check with the ground truth on whether a state belongs to the forbidden state set or to the goal state set.
Test Domains. We tested our method and baseline on five domains. Most of these are standard benchmark tasks taken from the SimpleRL library [Abel, 2019]. Since all examples were variations of a basic grid-world domain, we considered four different sizes and five random instantiations of each grid size (obtained by randomizing the initial state, goal state, forbidden states, and location of objects). For each of the tasks, the expectation set consists of reaching the goal state and avoiding some random states in the environment. The human models were generated by modifying the original task slightly. The walkway involves a simple grid world where the robot can use a movable walkway to reach certain states easily, but the human is unaware of it. Obstacles involve the robot navigating to a goal while avoiding obstacles (the human model includes incorrect information about the obstacles). Four rooms [Sutton et al., 1999] involves the robot navigating through connected rooms, but in the human model, the use of certain doors may not be allowed. In Puddle, the robot needs to travel to the goal while avoiding certain terrain types, while the human model may be wrong about the location of various terrain elements. Finally, in maze, the robot needs to navigate a maze, while the human model may have incorrect information about what paths are available.
Evaluation. All the baselines were run with a time-bound of 30 minutes per problem. All experiments were run on AlmaLinux 8.9 with 32GB RAM and 16 Intel(R) Xeon(R) 2.60GHz CPUs. We used CPLEX [Bliek1ú et al., 2014] as our LP solver (no-cost edition) [^4] The code for our experiments can be found at https://github.com/Malek-Mechergui/codeMDP. First, we found that the MMRQ-k method could not solve any of the instances. This was because the runtime was dominated by the time needed to calculate the exponential number of dominant policies. This shows how computationally expensive methods from the first group are. On the other hand, our method is much faster, and even in the largest grids, it took less than five minutes (Table 1) and only required very few queries. Note that the maximum number of queries that could be raised in each case corresponds to the total state space. In each case considered here the number of queries raised was substantially smaller than the state space size. Thus showing the effectiveness of our method compared to existing query methods.
@@ 6 Evaluation
Table 1: For our method, the table reports the number of queries raised and the time taken by our method. For IRD, it shows the number of expectations violated by the generated policy and the time taken. Note that our method is guaranteed not to choose a policy that results in violated expectations.